# Embodied AI in Urban Spaces

### v1

@@ A.1 Static Scene Samples
Street blocks. We design five typical street block categories – straight, curve, intersection, T-junction, and roundabout. In the simulator, to form a large map with several blocks, we can sample the category, number, order, lane number, and other related parameters of the blocks. We use the algorithm Block Incremental Generation (BIG) proposed in MetaDrive [44] to generate the target road network defined by users. Figure 8 provides demonstrations of generated street maps composed of different numbers of blocks.
@@ A.3 Static Asset Samples
We provide 10,000 high-quality static object assets. The roadside infrastructure is divided into three categories: 1) Standard infrastructure, including poles, trees, and signs, is placed at regular intervals along the road. 2) Non-standard infrastructure, such as buildings, bonsai, and trash bins, is placed randomly within designated zones. 3) Clutter, such as drink cans, bags, and bicycles, is scattered randomly across all functional zones. Figure 14 15 and 16 show examples of these three categories respectively.
@@ A.4 Dynamic Asset Samples
Human assets. MetaUrban provides 1,100 rigged 3D human models, sampled from 68 garments, 32 hairs, 13 beards, 46 accessories, and 1,038 cloth and skin textures from SynBody [81] dataset. Figure 17 shows randomly sampled humans, which have large variations.
@@ B.1 Layout Generation
Ground plan. As shown in Figure 21 (Top), we define 4 functional zones   and 6 geometric zones   for sampling the type of sidewalks and choosing the distribution of parameters for each sidewalk component. As shown in Figure 21 (Bottom), we construct 7 typical templates for sidewalks; each type of them has its unique distribution of geometric zones. To match the distribution with the real world, we set the distribution of the zone width to a uniform distribution for each geometric zone; the maximum and minimum values of the uniform distribution are set according to the Global Street Design Guide [31] provided by the Global Designing Cities Initiative.
@@
Object placement. Figure 22 illustrates the iterative process of placing objects in the scene. First, we convert the polygon of the geometric zone of the sidewalk into rectangles. We will place objects on each geometric zone independently. At each step of placing on the specific zone, we can obtain rectangles that are not occupied. Then we check from the starting region to the ending region for the current retrieved object class. We place it if possible, or we start to place the next class. In the simulator, we use rectangle bounding boxes to represent all objects physically to adopt this object placement method.
Figure 22: Iterative object placement. Given a ground plan, we first convert geometric zones into rectangles. Then, we place objects on zones one by one. In the placing of each zone, we attempt to put sampled classes of objects in the rectangles one by one.
@@ B.2 Object Retrieval
We first leverage off-the-shelf scene understanding datasets – Mapillary Vistas [52] and CityScape [10]. Using the provided annotation polygon, we find the overlapping object with the sidewalk and get a list of 90 objects that are with high frequency to be put in the urban space (such as “tree” and “bench”). However, the number of objects is limited because of the closed-set definitions in the image datasets. To get broader object distribution from the real world, we introduce two open-set sources – worldwide Google Street data and urban planning description data.
For the Google Street data, we collect 25,000 urban space images from 50 countries across six continents. The selection of image locations was performed by randomly sampling points along the major roads of cities using OpenStreetMap’s [54] road network. Image orientation was determined based on road gradient to enhance the relevance of captured scenes. For object detection in these images, we initially employed GPT-4o [53] to generate a list of candidate objects. This was followed by the application of Grounded-Dino[47] to obtain bounding boxes for these objects. We refined these boxes using non-maximum suppression (NMS) to ensure the accuracy of object identification.
@@
Further refinement was achieved through the use of the Grounded-SAM model [66], an open-set segmentation approach, which filtered the bounding boxes to identify objects specifically located in public urban spaces. A key part of our method involves determining overlaps between identified objects and sidewalks. For each object detected, we calculate its spatial intersection with sidewalk regions derived from the datasets. This overlap analysis helps in curating a list of objects that are relevant to urban public spaces.
To address the diverse descriptions generated by GPT-4o [53] and ensure semantic uniformity, we cluster the embeddings of descriptions using DBSCAN [21], which result in 1,075 distinct object clusters with unique descriptors, such as "a gray trash bin" and "potted cactus". We use “all-mpnet-base-v2” model from SentenceTransformers [65] to embed each description.
@@
Open-vocabulary search. To effectively retrieve digital assets corresponding to the object description pool, we developed a robust pipeline utilizing the Objaverse [17] and ObjaverseXL [15] repositories, known for their extensive digital assets. The process begins with the extraction of digital assets using a multi-threaded approach for further processing. Each downloaded asset is then rendered into 20 distinct images, capturing various angles to provide a comprehensive visual representation. Following [49, 48], viewpoints with higher quality are used for the calculation of visual feature embedding.
For the matching process, we leverage the BLIP2 [42] model, a pre-trained feature extractor, to align visual data with our textual descriptions. This involves processing the images to extract visual features and concurrently transforming textual descriptions into embeddings. These embeddings are compared using cosine similarity to determine the semantic correspondence between text and images, allowing us to identify and collect the digital assets that best match the descriptions.
@@
Object repository extension. MetaDrive provides an interface for including objects enabled by recent advances in 3D content generation, such as 3D object reconstruction [46, 34] and generation [60, 9]. Thus, one can easily further extend the object repository with generated contents. Also, this function can work together with scene customization (Section B.4) to get customized scenes with specific objects.
@@ B.3 Cohabitant Populating
Movements. We include 3 daily movements – idle, walking, and running, as well as 2,311 unique movements from the BEDLAM [6] dataset. All of the motion sequences are trimmed and checked by designers one by one to ensure their quality. With the same skeletal binding, all of the unique movements can be transferred to all of the 3D human models directly. Thus, we can get 1,100 $\times$ 2,311 numbers of human-motion pairs.
Trajectories. We harness ORCA [77] and Push and Rotate (P&R) algorithm [11] to get the trajectories of all dynamic agents. First, we build the 0-1 mask that indicates whether the grid is a walkable region or not. Then, we sample the start and ending points for each agent randomly, followed by generating their 2D trajectories by using the model of ORCA [77] and P&R [11]. The trajectory plan process is efficient, running within 5s for 100 agents on a Core i9 CPU processor. Vehicles will also be added in dynamic scenes. All traffic vehicles will follow IDM policies, as MetaDrive [44] does.
@@ B.6 Simulator Comparison
We will compare MetaUrban with other simulators below in Table 2, using the scale, sensor, and feature dimensions. For the scale, MetaUrban can generate infinite scenes with a procedural generation pipeline. It provides the largest number of humans (1,100) and movements (2,314) among all simulation environments. For objects, so far, we have provided 10,000. Compared to other simulators, all of the objects from MetaUrban are urban-specific. Also, we provide an interface to extend object data to any size easily with recent advances in 3D content generation (Section B.2). For the sensor, MetaUrban provides RGBD, semantic, and lidar. For the feature, different from other simulators, MetaUrban provides real-world distribution of the object’s categories and uses a more sophisticated path plan algorithm to get the natural agent’s trajectories. It also provides flexible user interfaces – mouse, keyboard, joystick, and racing wheel, which vastly ease the collection of human expert demonstration data. MetaUrban uses PyBullet as its physical engine and Panda3D [27] for rendering.
@@
HuNavSim [59] 5 5 6 Social Force Gazebo Social
SEAN 2.0 [76] 3 34 $<$100 1 Manual Social Force Unity Social
SocNavBench [5] 4 Social
SUMO [38] $\infty$ Driving
CARLA [19] 15 66,599 49 1 Manual Rule-based Keyboard, Joystick Unreal4 Driving
MetaDrive [44] $\infty$ 5 1 1 Manual Rule-based PyBullet Driving
AI2-THOR [35] 120 609 Manual Mouse Unity Indoor
ThreeDWorld [24] 15 200 Manual VR Flex Indoor
iGibson 2.0 [40] 15 1,217 Manual Mouse, VR PyBullet Indoor
ProcTHOR [18] $\infty$ 1,547 Manual Unity Indoor
OmniGibson [41] 306 5,215 Manual PhysX Indoor
Habitat 3.0 [63] 211 18,656 12 3 Manual Rule-based Mouse, Keyboard, VR Bullet Indoor
@@
Appendix C Experiment Details
@@ C.1 PointNav Experiments
We use the same continuous action space as MetaDrive [44], which is a 2-dimensional vector that normalized to $[-1.0,1.0]$ indicating the acceleration and steering rate of the agent. Considering that the dynamics of a delivery robot is different from a vehicle, we change some core parameters like maximum velocity, maximum acceleration, maximum steering rate and so on.
@@
For PointNav, an episode is considered successful if the agent issues the DONE action, defined as completing 95% of the set route within 1,000 maximum steps. The agent is evaluated using the Success Rate (SR) and Success weighted by Path Length (SPL) [2, 4] metrics, which measure the effectiveness and efficiency of the path taken by the agent. Additionally, to measure the safety performance of the trained policy, we define the cost function by two events, i.e., crashing with objects on the sidewalk or buildings in the building zone. $+1$ cost is given once those events occur.
@@ Methods.
Reinforcement learning. In the realm of Reinforcement Learning, we use the Proximal Policy Optimization (PPO) [68] for evaluation. PPO is a widely adopted and effective method that strikes a balance between sample complexity and ease of tuning, and it is easy to scale as it adopts parallel and distributed training well. The agent in this setting is trained to maximize the reward, which we carefully design to encapsulate the desired behavior of the agent in the MetaUrban environment. The specifics of the reward structure will be discussed in the subsequent paragraph. We train the PPO using the same set of hyperparameters with 128 parallel environments, which occupy 128 processes. The total training time is 12 hours, and 5M environment steps for PointNav on a single Nvidia A5000 GPU. The detailed hyperparameters are provided in Table 3.
@@
Safe reinforcement learning. As driving in urban spaces is a safety-critical application, it is important to evaluate Safe Reinforcement Learning (SafeRL) algorithms. In the domain of SafeRL, we utilize two approaches: PPO with a Lagrangian constraint (PPO-Lag) [64] and PPO with modeling of Early Terminated Markov Decision Processes (PPO-ET) [72]. Both methods aim to ensure that the learned policies adhere to specific safety constraints while optimizing the reward. PPO-Lag incorporates a Lagrangian term into the objective function to enforce the constraints, while PPO-ET changes the modeling of the Constrained Markov Decision Process (CMDP) to a new unconstrained MDP, the optimal policy that coincidences with the original CMDP.
For PPO-Lag [64], it considers the learning objectivate as Equation 1 rather than adding negative cost as rewards.
@@
The rule for PPO-ET [72] is to stop when the constraint cost exceeds a given value, which can be easily implemented in practice.
We implement both of these SafeRL methods based on OmniSafe [33]. We train both of them with 50 parallel environments and the training takes 12 hours for PointNav on a single Nvidia A5000 GPU. The detailed hyperparameters are provided in Table 3.
Offline reinforcement learning. For Offline Reinforcement Learning, we employ two prominent methods: Implicit Q-Learning (IQL) [36] and Twin Delayed Deep Deterministic Policy Gradient with Behavior Cloning (TD3+BC) [23]. We create the dataset for PointNav by combining $20\%$ human demonstrations with $80\%$ demonstrations from a well-trained PPO policy, consisting of 30,000 samples with approximately 60% success rate. The training is purely offline and takes around 2 hours on a single Nvidia A5000 GPU for 100 epochs. The detailed hyperparameters for IQL and TD3+BC are provided in Table 4 and 5, respectively.
@@
Imitation learning. For Imitation Learning algorithms, we use the same high-quality mixed demonstration used in Offline Reinforcement Learning. In the Imitation Learning setting, the agent learns to mimic the behavior shown in the expert demonstration, and it is differentiated from Offline Reinforcement Learning in the sense that the agent does not have access to the rewards. We employ two well-established methods: Behavior Cloning (BC) [3] and Generative Adversarial Imitation Learning (GAIL) [30]. BC is a straightforward approach that trains the agent to directly match the actions of the expert given the observed states. GAIL, on the other hand, formulates the imitation learning problem as a two-player game between the agent and a discriminator, which tries to distinguish between the agent’s behavior and the expert’s demonstrations. The detailed hyperparameters for IQL and TD3+BC are provided in table 6 and 7, respectively.
@@
  • Crash reward $R_{crash}$: a dense negative reward defined as $-1(c_{t})$, wherein the $c_{t}$ denotes the collision between agents and any other objects at time $t$ and $1(\cdot)$ is the indicator function. It’s notable we do not use the termination strategy for collision as in MetaDrive [44]. We set the weight of $R_{crash}$ as $c_{4}=1.0$.

### v2

@@ A.1 Static Scene Samples
Street blocks. We design five typical street block categories – straight, curve, intersection, T-junction, and roundabout. In the simulator, to form a large map with several blocks, we can sample the category, number, order, lane number, and other related parameters of the blocks. We use the algorithm Block Incremental Generation (BIG) proposed in MetaDrive (Li et al., 2022b) to generate the target road network defined by users. Figure 7 provides demonstrations of generated street maps composed of different numbers of blocks.
@@ A.3 Static Asset Samples
We provide 10,000 high-quality static object assets. The roadside infrastructure is divided into three categories: 1) Standard infrastructure, including poles, trees, and signs, is placed at regular intervals along the road. 2) Non-standard infrastructure, such as buildings, bonsai, and trash bins, is placed randomly within designated zones. 3) Clutter, such as drink cans, bags, and bicycles, is scattered randomly across all functional zones. Figure 13 14 and 15 show examples of these three categories respectively.
@@ A.4 Dynamic Asset Samples
Human assets. MetaUrban provides 1,100 rigged 3D human models, sampled from 68 garments, 32 hairs, 13 beards, 46 accessories, and 1,038 cloth and skin textures from SynBody Yang et al. (2023) dataset. Figure 16 shows randomly sampled humans, which have large variations.
@@ B.1 Layout Generation
Ground plan. As shown in Figure 20 (Top), we define 4 functional zones   and 6 geometric zones   for sampling the type of sidewalks and choosing the distribution of parameters for each sidewalk component. As shown in Figure 20 (Bottom), we construct 7 typical templates for sidewalks; each type of them has its unique distribution of geometric zones. To match the distribution with the real world, we set the distribution of the zone width to a uniform distribution for each geometric zone; the maximum and minimum values of the uniform distribution are set according to the Global Street Design Guide (Initiative & of City Transportation Officials, 2016) provided by the Global Designing Cities Initiative.
@@
Object placement. Figure 21 illustrates the iterative process of placing objects in the scene. First, we convert the polygon of the geometric zone of the sidewalk into rectangles. We will place objects on each functional zone or geometric zone independently. At each iteration of generating on the specific zone, we can obtain rectangles that are not occupied. Then we check from the starting region to the ending region for the current retrieved object class. We place it if possible, or we start to place the next class. In the simulator, we use rectangle bounding boxes to represent all objects physically to adopt this object placement method.
Figure 21: Iterative object placement.
@@ B.2 Object Retrieval
We first leverage off-the-shelf scene understanding datasets – Mapillary Vistas (Neuhold et al., 2017) and CityScape (Cordts et al., 2016). Using the provided annotation polygon, we find the overlapping object with the sidewalk and get a list of 90 objects that are with high frequency t1o be put in the urban space (such as “tree” and “bench”). However, the number of objects is limited because of the closed-set definitions in the image datasets. To get broader object distribution from the real world, we introduce two open-set sources – worldwide Google Street data and urban planning description data.
For the Google Street data, we collect 25,000 urban space images from 50 countries across six continents. The selection of image locations was performed by randomly sampling points along the major roads of cities using OpenStreetMap’s (OpenStreetMap contributors, 2017) road network. Image orientation was determined based on road gradient to enhance the relevance of captured scenes. For object detection in these images, we initially employed GPT-4o (OpenAI, 2024) to generate a list of candidate objects. This was followed by the application of Grounded-Dino (Liu et al., 2023c) to obtain bounding boxes for these objects. We refined these boxes using non-maximum suppression (NMS) to ensure the accuracy of object identification.
@@
Further refinement was achieved through the use of the Grounded-SAM model (Ren et al., 2024), an open-set segmentation approach, which filtered the bounding boxes to identify objects specifically located in public urban spaces. A key part of our method involves determining overlaps between identified objects and sidewalks. For each object detected, we calculate its spatial intersection with sidewalk regions derived from the datasets. This overlap analysis helps in curating a list of objects that are relevant to public urban spaces.
To address the diverse descriptions generated by GPT-4o (OpenAI, 2024) and ensure semantic uniformity, we cluster the embeddings of descriptions using DBSCAN (Ester et al., 1996), which result in 1,075 distinct object clusters with unique descriptors, such as "a gray trash bin" and "potted cactus". We use “all-mpnet-base-v2” model from SentenceTransformers (Reimers & Gurevych, 2019) to embed each description.
@@
Open-vocabulary search. To effectively retrieve digital assets corresponding to the object description pool, we developed a robust pipeline utilizing the Objaverse (Deitke et al., 2023c) and ObjaverseXL (Deitke et al., 2023a) repositories, known for their extensive digital assets. The process begins with the extraction of digital assets using a multi-threaded approach for further processing. Each downloaded asset is then rendered into 20 distinct images, capturing various angles to provide a comprehensive visual representation. Following (Luo et al., 2023; 2024), viewpoints with higher quality are used for the calculation of visual feature embedding.
For the matching process, we leverage the BLIP2 (Li et al., 2023) model, a pre-trained feature extractor, to align visual data with our textual descriptions. This involves processing the images to extract visual features and concurrently transforming textual descriptions into embeddings. These embeddings are compared using cosine similarity to determine the semantic correspondence between text and images, allowing us to identify and collect the digital assets that best match the descriptions.
@@
Object repository extension. MetaDrive provides an interface for including objects enabled by recent advances in 3D content generation, such as 3D object reconstruction (Liu et al., 2023b; Kerbl et al., 2023) and generation (Poole et al., 2023; Chen et al., 2023). Thus, one can easily further extend the object repository with generated contents. Also, this function can work together with scene customization (Section B.4) to get customized scenes with specific objects.
@@ B.3 Cohabitant Populating
Movements. We include 3 daily movements – idle, walking, and running, as well as 2,311 unique movements from the BEDLAM (Black et al., 2023) dataset. All of the motion sequences are trimmed and checked by designers one by one to ensure their quality. With the same skeletal binding, all of the unique movements can be transferred to all of the 3D human models directly. Thus, we can get 1,100 $\times$ 2,311 numbers of human-motion pairs.
Trajectories. We harness ORCA (Van Den Berg et al., 2011) and Push and Rotate (P&R) algorithm (De Wilde et al., 2014) to get the trajectories of all dynamic agents. First, we build the 0-1 mask that indicates whether the grid is a walkable region or not. Then, we sample the start and ending points for each agent randomly, followed by generating their 2D trajectories by using the model of ORCA (Van Den Berg et al., 2011) and P&R (De Wilde et al., 2014). The trajectory plan process is efficient, running within 5s for 100 agents on a Core i9 CPU processor. Vehicles will also be added in dynamic scenes. All traffic vehicles will follow IDM policies, as MetaDrive (Li et al., 2022b) does.
@@ B.6 Simulator Comparison
We will compare MetaUrban with other simulators below in Table 3, through the scale, sensor, and feature dimensions. For the scale, MetaUrban can generate infinite scenes with a procedural generation pipeline. It provides the largest number of humans (1,100) and movements (2,314) among all simulation environments. For objects, so far, we have provided 10,000. Compared to other simulators, all of the objects from MetaUrban are urban-specific. Also, we provide an interface to extend object data to any size easily with recent advances in 3D content generation (Section B.2). For the sensor, MetaUrban provides RGBD, semantic, and lidar. For the feature, different from other simulators, MetaUrban provides real-world distribution of the object’s categories and uses a more sophisticated path plan algorithm to get the natural agent’s trajectories. It also provides flexible user interfaces – mouse, keyboard, joystick, and racing wheel, which vastly ease the collection of human expert demonstration data. MetaUrban uses PyBullet as its physical engine, which is open-source and highly accurate in physics simulation, providing a cost-effective and flexible solution for future developments. MetaUrban uses Panda3D (Goslin & Mine, 2004) for rendering, which is a lightweight, open-source game engine with seamless Python integration, providing a flexible and accessible development environment.
@@
HuNavSim (Pérez-Higueras et al., 2023) 5 5 6 Social Force Gazebo Social
SEAN 2.0 (Tsoi et al., 2022) 3 34 <100 1 Manual Social Force Unity Social
SocNavBench (Biswas et al., 2022) 4 Social
SUMO (Krajzewicz et al., 2002) $\infty$ Driving
CARLA (Dosovitskiy et al., 2017) 15 66,599 49 1 Manual Rule-based Keyboard, Joystick Unreal4 Driving
MetaDrive (Li et al., 2022b) $\infty$ 5 1 1 Manual Rule-based PyBullet Driving
AI2-THOR (Kolve et al., 2017) 120 609 Manual Mouse Unity Indoor
ThreeDWorld (Gan et al., 2021) 15 200 Manual VR Flex Indoor
iGibson 2.0 (Li et al., 2021) 15 1,217 Manual Mouse, VR PyBullet Indoor
ProcTHOR (Deitke et al., 2022b) $\infty$ 1,547 Manual Unity Indoor
OmniGibson (Li et al., 2024) 306 5,215 Manual PhysX Indoor
Habitat 3.0 (Puig et al., 2023b) 211 18,656 12 3 Manual Rule-based Mouse, Keyboard, VR Bullet Indoor
@@
Appendix C MetaUrban-12K Dataset
Data.
[…]
Statistics.
[…]
Figure 25: MetaUrban-12K statistics. (Left) Distribution of object numbers in the scene. (Middle) Distribution of areas occupied by objects in the scene. (Right) Distribution of episode length in the scene.
Appendix D Experiment Details
@@ D.1 PointNav Experiments
We use the same continuous action space as MetaDrive (Li et al., 2022b), which is a 2-dimensional vector that normalized to $[-1.0,1.0]$ indicating the acceleration and steering rate of the agent. Considering that the dynamics of a delivery bot is different from a vehicle, we change some core parameters like maximum velocity, maximum acceleration, maximum steering rate.
@@
For PointNav, an episode is considered successful if the agent issues the DONE action, defined as completing 95% of the set route within 1,000 maximum steps. The agent is evaluated using the Success Rate (SR) and Success weighted by Path Length (SPL) (Anderson et al., 2018; Batra et al., 2020) metrics, which measure the effectiveness and efficiency of the path taken by the agent. Additionally, to measure the safety performance of the trained policy, we define the cost function by two events, i.e., crashing with objects on the sidewalk or buildings in the building zone. $+1$ cost is given once those events occur.
@@ Methods.
Reinforcement learning. In the realm of Reinforcement Learning, we use the Proximal Policy Optimization (PPO) (Schulman et al., 2017) for evaluation. PPO is a widely adopted and effective method that strikes a balance between sample complexity and ease of tuning, and it is easy to scale as it adopts parallel and distributed training well. The agent in this setting is trained to maximize the reward, which we carefully design to encapsulate the desired behavior of the agent in the MetaUrban environment. The specifics of the reward structure will be discussed in the subsequent paragraph. We train the PPO using the same set of hyperparameters with 128 parallel environments, which occupy 128 processes. The total training time is 12 hours, and 5M environment steps for PointNav on a single Nvidia A5000 GPU. The detailed hyperparameters are provided in Table 4.
@@
Safe reinforcement learning. As driving in urban spaces is a safety-critical application, it is important to evaluate Safe Reinforcement Learning (SafeRL) algorithms. In the domain of SafeRL, we utilize two approaches: PPO with a Lagrangian constraint (PPO-Lag) (Ray et al., 2019) and PPO with modeling of Early Terminated Markov Decision Processes (PPO-ET) (Sun et al., 2021). Both methods aim to ensure that the learned policies adhere to specific safety constraints while optimizing the reward. PPO-Lag incorporates a Lagrangian term into the objective function to enforce the constraints, while PPO-ET changes the modeling of the Constrained Markov Decision Process (CMDP) to a new unconstrained MDP, the optimal policy that coincidences with the original CMDP.
For PPO-Lag (Ray et al., 2019), it considers the learning objectivate as Equation 1 rather than adding negative cost as rewards.
@@
The rule for PPO-ET (Sun et al., 2021) is to stop when the constraint cost exceeds a given value, which can be easily implemented in practice.
We implement both of these SafeRL methods based on OmniSafe (Ji et al., 2023). We train both of them with 50 parallel environments and the training takes 12 hours for PointNav on a single Nvidia A5000 GPU. The detailed hyperparameters are provided in Table 4.
Offline reinforcement learning. For Offline Reinforcement Learning, we employ two prominent methods: Implicit Q-Learning (IQL) (Kostrikov et al., 2021) and Twin Delayed Deep Deterministic Policy Gradient with Behavior Cloning (TD3+BC) (Fujimoto & Gu, 2021). We create the dataset for PointNav by combining $20\%$ human demonstrations with $80\%$ demonstrations from a well-trained PPO policy, consisting of 30,000 samples with approximately 60% success rate. The training is purely offline and takes around 2 hours on a single Nvidia A5000 GPU for 100 epochs. The detailed hyperparameters for IQL and TD3+BC are provided in Table 5 and 6, respectively.
@@
Imitation learning. For Imitation Learning algorithms, we use the same high-quality mixed demonstration used in Offline Reinforcement Learning. In the Imitation Learning setting, the agent learns to mimic the behavior shown in the expert demonstration, and it is differentiated from Offline Reinforcement Learning in the sense that the agent does not have access to the rewards. We employ two well-established methods: Behavior Cloning (BC) (Bain & Sammut, 1995) and Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016). BC is a straightforward approach that trains the agent to directly match the actions of the expert given the observed states. GAIL, on the other hand, formulates the imitation learning problem as a two-player game between the agent and a discriminator, which tries to distinguish between the agent’s behavior and the expert’s demonstrations. The detailed hyperparameters for IQL and TD3+BC are provided in table 7 and 8, respectively.
@@
  • Crash reward $R_{crash}$: a dense negative reward defined as $-1(c_{t})$, wherein the $c_{t}$ denotes the collision between agents and any other objects at time $t$ and $1(\cdot)$ is the indicator function. It’s notable we do not use the termination strategy for collision as in MetaDrive (Li et al., 2022b). We set the weight of $R_{crash}$ as $c_{4}=1.0$.